[Hallc_running] [New Logentry] Hall C RC update, Thurs 8/30

mack at jlab.org mack at jlab.org
Thu Aug 30 09:05:02 EDT 2018

Logentry Text:
Due to NL vacuum problems*, there will be no beam today. Since the source of the problem is still speculative, we'll revisit the situation after Friday's 8AM meeting. For now, we'll keep the TO watching a cold target, but the following SL shifts are cancelled: day Thurs (Burcu), evening Thurs (Rolf), and owl Friday (Holly). 

We have asked to take Hall C to controlled access. This could take a while since the NL work has priority. 

Let Walter and me know of significant work plans for Hall C. The only thing I know of is Lassiter/Kellner checking the minimum opening angle of the spectrometers. Dave G was interested in having part of the beamline surveyed but nothing has been arranged yet. Please think of useful things to do to improve our readiness for data taking, but don't make changes that can't be restored in an hour. 

*Yesterday, when the MCC tried to go CW, the vacuum deteriorated in NL21/22/23 . It was problematic even before that. They can't keep the valves open at reasonable setpoints. The working hypothesis is there's a ball of air ice in that region from a leak. This is the same region where they have had to steer to one side to avoid an apparent restriction. Arne noted that if the speculative hypothesis holds up, the time constant for repairs is set by how long it takes to warm up a massive cryomodule, potentially1-2 weeks. 

We will have our usual Hall C RC meeting today at 4:15 pm: https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_runplan_fall_2018


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3592935
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180830/fea298ee/attachment.html>

More information about the Hallc_running mailing list